A magnetic field is produced inside a cubic chamber of volume A and the electromagnetic waves propagate at the speed of light c. If some other particle is present in this chamber, list the four conditions under which this particle of mass m will experience a force-free "life".


1
Expert's answer
2021-08-26T14:48:20-0400

Given that mass of the particles inside the chamber be m and the volume of the chamber be A.

Particle will experience a force free life if

  1. "mg=qE"
  2. "\frac{ mv^2}{R^2} = qVB"
  3. If electric and magnetic field are perpendicular in such a way that "qvB+qE=mg"
  4. If electric field and magnetic field arrange in such a way that "qvB=mg+qE" or "qE= qvB+mg"